40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 21 19:28:03 2021, Anchal, Update, PSL, PSL/IFO recovery 
    Reply  Wed Apr 21 19:43:20 2021, Koji, Update, PSL, New HEPA speed control P_20210421_193637.jpgP_20210421_193627.jpg
       Reply  Thu Apr 22 14:41:55 2021, Chub, Update, PSL, New HEPA speed control 40M_PSL_HEPA.jpg
          Reply  Fri Apr 23 18:00:02 2021, gautam, Update, PSL, HEPA speed lowered HEPAdiag.pdf
             Reply  Fri Apr 23 19:26:58 2021, Koji, Update, PSL, HEPA speed lowered 
             Reply  Fri May 14 17:45:05 2021, rana, Update, PSL, HEPA speed raised 
                Reply  Tue May 18 00:52:38 2021, rana, Update, PSL, HEPA speed raised Untitled.png
                   Reply  Tue May 18 20:26:11 2021, rana, Update, PSL, HEPA speed raised 
    Reply  Thu Apr 22 15:15:26 2021, gautam, Update, PSL, PMC transmission 
       Reply  Thu Apr 22 15:34:54 2021, Anchal, Update, PSL, PMC transmission 
Message ID: 16077     Entry time: Thu Apr 22 15:34:54 2021     In reply to: 16076
Author: Anchal 
Type: Update 
Category: PSL 
Subject: PMC transmission 

Koji mentioned that the mode of the laser is different for lower diode currents. So that might be the reason why we got less transmission at the low input power but more afterward.

ELOG V3.1.3-